# Spectral gap bounds for reversible hybrid Gibbs chains

### v3

@@ Abstract
Keywords: convergence rate, data augmentation, Metropolis-within-Gibbs, MCMC, slice sampler.
@@ 1 Introduction
Despite the pervasive popularity and extensive application of these hybrid algorithms, significant gaps persist in our understanding of their convergence properties. Building on the foundational works of Roberts and Rosenthal [64] and Andrieu et al. [3], this article aims to bridge the gap by establishing a connection between the convergence rates of exact and hybrid Gibbs chains in some important scenarios. Intuitively, if the Markovian steps in a hybrid Gibbs sampler approximate the intractable conditional distributions well, then the convergence rate of the hybrid Gibbs sampler should be similar to that of the corresponding exact Gibbs sampler. Our analysis confirms this intuition and makes this notion precise with some quantitative convergence rate inequalities. In particular, we provide bounds on the absolute spectral gap of the hybrid algorithm in terms of that of the exact algorithm and the quality of the approximation. Bounds involving Dirichlet form and asymptotic variance are also provided.
@@
  • Secondly, our theory can be used to generalize existing frameworks for analyzing Gibbs algorithms. In Section 6.2, we apply our method to analyze random-scan Gibbs samplers that update multiple coordinates at a time. In particular, we provide a quantitative relationship between two samplers that update, respectively, $\ell$ and $\ell^{\prime}$ coordinates at a time, in terms of their spectral gaps. This extends the main result (Theorem 2) in [63], which itself is a generalization of a breakthrough technique called “spectral independence” recently developed in theoretical computer science for analyzing the convergence properties of Gibbs algorithms [1, 21, 10, 28].
@@
We now review some relevant existing works and explain how our contribution relates to them. On finite state spaces, Liu [45] constructed a Metropolized random-scan Gibbs sampler and showed that it is asymptotically more efficient than the corresponding exact Gibbs sampler using Peskun’s ordering [58]. Analysis on general state spaces is more challenging. The convergence properties of hybrid Gibbs samplers were studied qualitatively by Roberts and Rosenthal [64], Fort et al. [29], and Qin and Jones [62]. Quantitative studies on the subject remain scarce. When the approximating Markov chains are uniformly ergodic with convergence rates bounded away from one, some quantitative convergence bounds are given in [65], [57], and [38]. For hybrid random-scan Gibbs algorithms, we develop a spectral gap bound under the weaker assumption that the approximating chains are geometrically ergodic with convergence rates bounded away from unity. This is our main contribution. We also develop a new spectral gap bound for hybrid data augmentation algorithms. For these algorithms, Andrieu et al. [3] gave results leading to a spectral gap bound that is similar to our bound for hybrid random-scan Gibbs algorithms in form. Building on their work, we provide an alternative form of bound for hybrid data augmentation algorithms. The proofs of our results utilize the idea of Markov decomposition [15, 51]. Some of the linear algebraic techniques that we employ are adopted from [3] and [41]. Following the initial release of our work, Ascolani et al. [7] derived bounds for the $s$-conductance of a class of hybrid random-scan Gibbs algorithms, exhibiting some similarities to our spectral gap bounds.
@@
The Appendix includes technical results and proofs, along with some additional contributions. In Appendix E, we derive a spectral gap bound for a type of hybrid proximal sampler using techniques from [3], rather than the methods from Sections 4 and 5, which may be of interest to practitioners. In Appendix F, we derive an alternative spectral gap bound for hybrid random-scan Gibbs algorithms using arguments from [64], and compare it to the bound derived in Section 4. While this alternative bound performs less favorably in high-dimensional settings, its derivation based on [64] may still offer valuable insights.
@@ 2.2 Summary of results
For hybrid data augmentation chains, the following convergence bound, which is also of the form (4), was derived by Andrieu et al. [3].
@@
If the Markov operator $Q_{1,z}$ is positive semi-definite for each $z$, then the right-hand-side can be tightened to $1-\|S\|_{M_{1}}$. In fact, [3] derived a more general form of the bound. The general form involves an intractable quantity, and bounding the intractable quantity in an obvious manner yields (5).
@@ 3 Preliminaries
For any positive integer $t$, let $K^{t}$ be the $t$-step Mtk. It is well-known that $K$ can be seen as a linear operator on $L_{0}^{2}(\omega)$: for $f\in L_{0}^{2}(\omega)$,
@@ 5 Data augmentation chain
Building on the work of Andrieu et al. [3], this section will present some new results regarding the relation between exact and hybrid data augmentation samplers.
@@
[3, Theorem 25, Remark 26] Suppose that there exist constants $c_{1},c_{2}\in[0,2]$ such that
@@ 6.1.1 Bayesian regression with a spike-and-slab prior
Consider a random-scan Gibbs sampler targeting $\Pi$. There are three full conditional distributions. We will let $\Pi_{1,(z,q)}(\cdot)$ be the conditional distribution of $\beta$ given $(z,q)$, and define $\Pi_{2,(\beta,q)}$ and $\Pi_{3,(\beta,z)}$ analogously. The second and third full conditional distributions may be sampled from efficiently using rejection sampling, since they are either univariate or a product of univariate distributions. On the other hand, it may be costly to sample exactly from $\Pi_{1,(z,q)}(\cdot)$. A random walk Metropolis-Hastings step may be used to approximate $\Pi_{1,(z,q)}(\cdot)$, as we describe below. Given the current state $\beta$, independently draw $\beta^{\prime}_{j}$, $j\in\{1,\dots,d\}$, from the $\mbox{N}(\beta_{j},\tau(z_{j})\sigma_{d}^{2})$ distribution, where $\sigma_{d}>0$. The proposed state $\beta^{\prime}$ is then accepted or rejected based on a Metropolis-Hastings step, ensuring that the transition law is reversible with respect to $\Pi_{1,(z,q)}(\cdot)$. Let $Q_{1,(z,q)}$ be the corresponding transition kernel. Let $T$ be the Mtk of an exact random-scan Gibbs sampler, i.e.,
@@
In cases where it is feasible but costly to implement the exact Gibbs sampler, $\mbox{var}_{\hat{T}}(f)/\mbox{var}_{T}(f)$ can be compared to the ratio of the average per-iteration costs of $T$ and $\hat{T}$ to determine the more efficient sampler. For illustration, suppose that $\ell(\beta)$ corresponds to a probit regression model based on $N_{d}$ iid observations, where the cost of evaluating $\ell(\beta)$ scales as $dN_{d}$. Then the cost of drawing from $Q_{1,(z,q)}(\beta,\cdot)$ is on the order of $dN_{d}$. On the other hand, some algorithms allow exact sampling from $\Pi_{1,(z,q)}$ [12, 26], such as Algorithm 1 in [26]. The per-iteration cost of this algorithm after preprocessing appears to scale as $dN_{d}+d^{2}$, plus the cost of sampling from an $N_{d}$-dimensional truncated normal distribution with $N_{d}$ linear constraints. Sampling from the truncated normal distributions can be performed using a rejection sampler, but the cost, denoted by $C(N_{d})$, may be prohibitively large for large $N_{d}$ [23, 12]. The costs of updating $z$ and $q$ are assumed to be negligible compared to the cost of drawing from $Q_{1,(z,q)}(\beta,\cdot)$ or $\Pi_{1,(z,q)}(\cdot)$. Thus, the ratio of the average per-iteration costs of $T$ and $\hat{T}$ is $\Omega(1+dN_{d}^{-1}+d^{-1}N_{d}^{-1}C(N_{d}))$. Comparing this to the asymptotic order of $\mbox{var}_{\hat{T}}(f)/\mbox{var}_{T}(f)$ in the previous paragraph shows that the Metropolis-within-Gibbs algorithm is more efficient if
$C(N_{d})\geq d^{2}N_{d}(N_{d}+d).$
@@ 6.1.2 A generic setting
      There exists $k\in{1,\dots,d}$ such that, for $i\leq k$ and $y\in\mathsf{X}_{-i}$, the log density $\xi_{i,y}$ is well conditioned in the following sense: there exist positive numbers $m_{i,y}$ and $L_{i,y}$ such that, for $z,u\in\mathbb{R}^{d_{i}}$,
@@ 6.3 A hybrid slice sampler
[55, Theorem 3.10, Proposition 3.13] Suppose that $v$ is continuously differentiable on the interval $(0,\|\tilde{m}\|_{\infty})$ with $v^{\prime}(z)<0$. Suppose further that, for some positive integer $k$, the function $z\mapsto zv^{\prime}(z)/v(z)^{1-1/k}$ is decreasing on $(0,\|\tilde{m}\|_{\infty})$. Then
@@
Suppose that, in Proposition 21, $\gamma(z)<1$ almost everywhere on $(0,\|\tilde{m}\|_{\infty})$, and that $\gamma(z)\leq\gamma_{0}$ when $z\in(0,\epsilon]$ for some $\gamma_{0}<1$ and $\epsilon>0$. Then, taking
@@
Assuming that $z\mapsto r_{z}/R_{z}$ is measurable in Proposition 23, we can take $\gamma(z)$ to be $1-(C_{1}/d^{2})(r_{z}/R_{z})^{2}$. By Proposition 22, in this case, to ensure that $\alpha_{t}$ converges to 0, it suffices to have $r_{z}/R_{z}$ being bounded away from 0 when $z\to 0$. That is, when $z\approx 0$, the level set $G_{z}$ is not excessively stretched and elongated.
@@ 7 Discussion
To obtain a non-trivial lower bound on $1-\|\hat{T}\|_{\Pi}$ using Corollary 4, it is imperative that $\|Q_{i,y}\|_{\Pi_{i,y}}\leq C$ for all values of $i$ and $y$, where $C$ is some constant that is strictly less than one. Relaxing this constraint emerges as a clear avenue for prospective research.
@@ Appendix E A proximal sampler
This Appendix contains an application of Proposition 7. Although the proposition has been discovered by Andrieu et al. [3], this application, which involves hybrid proximal samplers, appears new and may be interesting in its own right.
@@
In this case, the data augmentation chain associated with the Mtk $S$, where $S(y,\mathrm{d}y^{\prime})=\int_{\mathbb{R}^{d}}\Pi_{2,y}(\mathrm{d}z)\,\Pi_{1,z}(\mathrm{d}y^{\prime})$ is called a proximal sampler [43]. It can be particularly useful when $M_{1}$ is ill-conditioned, although later on in our analysis of its hybrid version, we focus on the case where $M_{1}$ is well-conditioned for tractability. This sampler was studied by Lee et al. [44] and Chen et al. [19]. In particular, the following holds.
@@
[19, Theorem 4] Suppose that $m>0$. Then $\|S\|_{M_{1}}\leq 1/(1+m\eta)$.
@@
Proposition 7 and Corollary 9 both provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{1,z}\|_{\Pi_{1,z}}$. In particular, combining the bound in Proposition 7 with Lemma 30 yields the following.
@@ Appendix F A bound from Roberts and Rosenthal (1997)
As mentioned in Section 1, [64] is a pioneering work on the convergence properties of hybrid Gibbs samplers. Roberts and Rosenthal [64] did not provide explicit bounds relating $\|T\|_{\Pi}$ to $\|\hat{T}\|_{\Pi}$, but it is possible to use their argument to obtain a quantitative relationship. In this section, we derive such a relation, and compare it to Corollary 4.
We first review the relevant result in [64]. For $i=1,\dots,n$, let $\Pi_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be such that $\Pi_{i}f(x)=\int_{\mathsf{X}}\Pi_{i,x_{-i}}(\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x_{-i}^{\prime})\,f(x^{\prime})$, and let $Q_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be such that $Q_{i}f(x)=\int_{\mathsf{X}}Q_{i,x_{-i}}(x_{i},\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x_{-i}^{\prime})\,f(x^{\prime})$. Then, $T$ and $\hat{T}$, as defined in (1) and (2), can be written as $T=\sum_{i=1}^{n}p_{i}\Pi_{i}$ and $\hat{T}=\sum_{i=1}^{n}p_{i}Q_{i}$, respectively.
@@
The quantity $C_{i,t}$ is slightly mysterious. Using the notation $f_{i,x_{-i}}(x_{i})=f(x)$ in the proof of Theorem 2, we may write it as follows:
@@
Assume, as in Corollary 4 that $\|Q_{i,y}\|_{\Pi_{i,y}}\leq C$ for some constant $C\in[0,1]$. Then
@@
Hence, (23) is looser than the lower bound in (12) from Corollary 4. For a rough quantitative comparison, fix $T$ and assume that $C=1-\varepsilon$ where $\varepsilon=o(1)$. Take $t=\lceil-\varepsilon^{-1}\log[(1-\|T\|_{\Pi})/2]+O(1)$. Then
@@
where $\Theta$ means the left and righ hand sides are of the same order. Then the lower bound in (23) is exponential in $-1/\varepsilon$, while the lower bound in (12) is proportional to $\varepsilon$. In the example in Section 6.1.2, $\varepsilon$ is inversely proportional to a polynomial of some dimension parameters. In that example, if $k=n$, $m_{i,y}/L_{i,y}$ is lower bounded by a positive constant $1/\kappa$, and $d_{i}=d$ for $i=1,\dots,n$, then one may take $\varepsilon=C_{0}/(\kappa d)$. In this case, (12) gives the bound $1-\|\hat{T}\|_{\Pi}\geq C_{0}/(\kappa d)(1-\|T\|_{\Pi})$, while (23) gives $1-\|\hat{T}\|_{\Pi}\geq c[(1-\|T\|_{\Pi})/2]^{(\log n)\kappa d/C_{0}}$ when $d$ is large, where $c$ is a positive constant. Evidently, the former is much tighter.
@@
Acknowledgment: We thank an Associate Editor and two Referees for helpful comments. An anonymous Referee provided valuable suggestions that helped streamline the proof of Theorem 2. We gratefully acknowledge the 2023 IRSA Conference at the University of Minnesota for catalyzing the collaborative efforts among them.
The first author was partially supported by the National Science Foundation through grant DMS-2112887. The third author was partially supported by the National Science Foundation through grant DMS-2210849 and an Adobe Data Science Research Award.

### v4

@@ Abstract
60J05,
convergence rate,
data augmentation,
Metropolis-within-Gibbs,
MCMC,
slice sampler,
keywords:
[class=MSC]
keywords:
[…]
, and
@@ 1 Introduction
Despite the pervasive popularity and extensive application of these hybrid algorithms, significant gaps persist in our understanding of their convergence properties. Building on the foundational works of Roberts and Rosenthal [64] and Andrieu, Lee and Vihola [2], this article aims to bridge the gap by establishing a connection between the convergence rates of exact and hybrid Gibbs chains in some important scenarios. Intuitively, if the Markovian steps in a hybrid Gibbs sampler approximate the intractable conditional distributions well, then the convergence rate of the hybrid Gibbs sampler should be similar to that of the corresponding exact Gibbs sampler. Our analysis confirms this intuition and makes this notion precise with some quantitative convergence rate inequalities. In particular, we provide bounds on the absolute spectral gap of the hybrid algorithm in terms of that of the exact algorithm and the quality of the approximation. Bounds involving Dirichlet form and asymptotic variance are also provided.
@@
  • Secondly, our theory can be used to generalize existing frameworks for analyzing Gibbs algorithms. In Section 6.2, we apply our method to analyze random-scan Gibbs samplers that update multiple coordinates at a time. In particular, we provide a quantitative relationship between two samplers that update, respectively, $\ell$ and $\ell^{\prime}$ coordinates at a time, in terms of their spectral gaps. This extends the main result (Theorem 2) in [63], which itself is a generalization of a breakthrough technique called ”spectral independence” recently developed in theoretical computer science for analyzing the convergence properties of Gibbs algorithms [1, 19, 10, 28].
@@
We now review some relevant existing works and explain how our contribution relates to them. On finite state spaces, Liu [45] constructed a Metropolized random-scan Gibbs sampler and showed that it is asymptotically more efficient than the corresponding exact Gibbs sampler using Peskun’s ordering [58]. Analysis on general state spaces is more challenging. The convergence properties of hybrid Gibbs samplers were studied qualitatively by Roberts and Rosenthal [64], Roberts and Rosenthal [65], Fort et al. [29], and Qin and Jones [62]. Quantitative studies on the subject remain scarce. When the approximating Markov chains are uniformly ergodic with convergence rates bounded away from one, some quantitative convergence bounds are given in [65], [57], and [38]. For hybrid random-scan Gibbs algorithms, we develop a spectral gap bound under the weaker assumption that the approximating chains are geometrically ergodic with convergence rates bounded away from unity. This is our main contribution.
We also develop a new spectral gap bound for hybrid data augmentation algorithms. For these algorithms, Andrieu, Lee and Vihola [2] gave results leading to a spectral gap bound that is similar to our bound for hybrid random-scan Gibbs algorithms in form. Building on their work, we provide an alternative form of bound for hybrid data augmentation algorithms.
The proofs of our results utilize the idea of Markov decomposition [15, 51]. Some of the linear algebraic techniques that we employ are adopted from [2] and [41]. Following the initial release of our work, Ascolani, Roberts and Zanella [7] derived bounds for the $s$-conductance of a class of hybrid random-scan Gibbs algorithms, exhibiting some similarities to our spectral gap bounds.
@@
The Appendix includes technical results and proofs. In a supplementary file, we derive a spectral gap bound for a type of hybrid proximal sampler using techniques from [2], rather than the methods from Sections 4 and 5, which may be of interest to practitioners. In the same file, we derive an alternative spectral gap bound for hybrid random-scan Gibbs algorithms using arguments from [64], and compare it to the bound derived in Section 4. While this alternative bound performs less favorably in high-dimensional settings, its derivation based on [64] may still offer valuable insights.
@@ 2.2 Summary of results
For hybrid data augmentation chains, the following convergence bound, which is also of the form (4), was derived by Andrieu, Lee and Vihola [2].
@@
If the Markov operator $Q_{1,z}$ is positive semi-definite for each $z$, then the right-hand-side can be tightened to $1-\|S\|_{M_{1}}$. In fact, Andrieu, Lee and Vihola [2] derived a more general form of the bound. The general form involves an intractable quantity, and bounding the intractable quantity in an obvious manner yields (5).
@@ 3 Preliminaries
For any positive integer $t$, let $K^{t}$ be the $t$-step Mtk, so that $K^{1}=K$ and $K^{t+1}(y,\cdot)=\int_{\mathsf{Y}}K(y,\mathrm{d}z)K^{t}(z,\cdot)$. It is well-known that $K$ can be seen as a linear operator on $L_{0}^{2}(\omega)$: for $f\in L_{0}^{2}(\omega)$,
@@ 5 Data augmentation chain
Building on the work of Andrieu, Lee and Vihola [2], this section will present some new results regarding the relation between exact and hybrid data augmentation samplers.
@@
[2, Theorem 25, Remark 26] Suppose that there exist constants $c_{1},c_{2}\in[0,2]$ such that
@@ 6.1.1 Bayesian regression with a spike-and-slab prior
Consider a random-scan Gibbs sampler targeting $\Pi$. There are three full conditional distributions. We will let $\Pi_{1,(z,q)}(\cdot)$ be the conditional distribution of $\beta$ given $(z,q)$, and define $\Pi_{2,(\beta,q)}$ and $\Pi_{3,(\beta,z)}$ analogously. The second and third full conditional distributions may be sampled from efficiently using rejection sampling, since they are either univariate or a product of univariate distributions. On the other hand, it may be costly to sample exactly from $\Pi_{1,(z,q)}(\cdot)$. A random walk Metropolis-Hastings step may be used to approximate $\Pi_{1,(z,q)}(\cdot)$, as we now describe:
  • Given the current state $\beta$, independently draw $\beta^{\prime}_{j}$, $j\in\{1,\dots,d\}$, from the $\mbox{N}(\beta_{j},\tau(z_{j})\sigma_{d}^{2})$ distribution, where $\sigma_{d}>0$.
  • The proposed state $\beta^{\prime}$ is then accepted or rejected based on a Metropolis-Hastings step, ensuring that the transition law is reversible with respect to $\Pi_{1,(z,q)}(\cdot)$.
Let $Q_{1,(z,q)}$ be the corresponding transition kernel. Let $T$ be the Mtk of an exact random-scan Gibbs sampler, i.e.,
@@
In cases where it is feasible but costly to implement the exact Gibbs sampler, the ratio $\mbox{var}_{\hat{T}}(f)/\mbox{var}_{T}(f)$ can be compared to the ratio of the average per-iteration costs of $T$ and $\hat{T}$ to determine the more efficient sampler. For illustration, suppose that $\ell(\beta)$ corresponds to a probit regression model based on $N_{d}$ iid observations, where the cost of evaluating $\ell(\beta)$ scales as $dN_{d}$. Then the cost of drawing from $Q_{1,(z,q)}(\beta,\cdot)$ is on the order of $dN_{d}$. On the other hand, some algorithms allow exact sampling from $\Pi_{1,(z,q)}$ [12, 26], such as Algorithm 1 in [26]. The per-iteration cost of this algorithm after preprocessing appears to scale as $dN_{d}+d^{2}$, plus the cost of sampling from an $N_{d}$-dimensional truncated normal distribution with $N_{d}$ linear constraints. Sampling from the truncated normal distributions can be performed using a rejection sampler, but the cost, denoted by $C(N_{d})$, may be prohibitively large for large $N_{d}$ [23, 12]. The costs of updating $z$ and $q$ are assumed to be negligible compared to the cost of drawing from $Q_{1,(z,q)}(\beta,\cdot)$ or $\Pi_{1,(z,q)}(\cdot)$. Thus, the ratio of the average per-iteration costs of $T$ and $\hat{T}$ is $\Omega(1+dN_{d}^{-1}+d^{-1}N_{d}^{-1}C(N_{d}))$. Comparing this to the asymptotic order of $\mbox{var}_{\hat{T}}(f)/\mbox{var}_{T}(f)$ in the previous paragraph shows that the Metropolis-within-Gibbs algorithm is more efficient if
$C(N_{d})\gg d^{2}N_{d}(N_{d}+d).$
@@ 6.1.2 A generic setting
      There exists $k\in\{1,\dots,d\}$ such that, for $i\leq k$ and $y\in\mathsf{X}_{-i}$, the log density $\xi_{i,y}$ is well conditioned in the following sense: there exist positive numbers $m_{i,y}$ and $L_{i,y}$ such that, for $z,u\in\mathbb{R}^{d_{i}}$,
@@ 6.3 A hybrid slice sampler
[55, Theorem 3.10, Proposition 3.13] Suppose that $v$ is continuously differentiable on the bounded interval $(0,\|\tilde{m}\|_{\infty})$ with $v^{\prime}(z)<0$. Suppose further that, for some positive integer $k$, the function $z\mapsto zv^{\prime}(z)/v(z)^{1-1/k}$ is decreasing on $(0,\|\tilde{m}\|_{\infty})$. Then
@@
Suppose that, in Proposition 21, $\gamma(z)<1$ almost everywhere on the bounded interval $(0,\|\tilde{m}\|_{\infty})$, and that $\gamma(z)\leq\gamma_{0}$ when $z\in(0,\epsilon]$ for some $\gamma_{0}<1$ and $\epsilon>0$. Then, taking
@@
Assuming that $z\mapsto r_{z}/R_{z}$ is measurable in Proposition 23, we can take $\gamma(z)=1-(C_{1}/d^{2})(r_{z}/R_{z})^{2}$. By Proposition 22, in this case, to ensure that $\alpha_{t}$ converges to 0, it suffices to have $r_{z}/R_{z}$ being bounded away from 0 when $z\to 0$. That is, when $z\approx 0$, the level set $G_{z}$ is not excessively stretched and elongated.
@@ 7 Discussion
To obtain a non-trivial lower bound on $1-\|\hat{T}\|_{\Pi}$ using Corollary 4, it is imperative that $\|Q_{i,y}\|_{\Pi_{i,y}}\leq C$ for all values of $i$ and $y$, where $C$ is some constant that is strictly less than one. Relaxing this constraint emerges as a clear avenue for prospective research. Techniques based on weak Poincaré inequalities or the $s$-conductance may be useful; see [4] and [7].
@@ Appendix D Proof of Proposition 22
{acks}
[Acknowledgments] We thank an Editor, an Associate Editor and two Referees for helpful comments. One Referee provided valuable suggestions that helped streamline the proof of Theorem 2. We gratefully acknowledge the 2023 IRSA Conference at the University of Minnesota for catalyzing the collaborative efforts among us. {funding} The first author was partially supported by the National Science Foundation through grant DMS-2112887. The third author was partially supported by the National Science Foundation through grant DMS-2210849 and an Adobe Data Science Research Award.
[…]
Supplement to ”Spectral gap bounds for reversible hybrid Gibbs chains”
This supplementary file contains two sections.
Section E provides an analysis of a Metropolis-within-proximal sampler based on Proposition 7 of the main text, taken from [2]. When the target distribution is well-conditioned, it is shown that the spectral gap of the hybrid proximal sampler is bounded from below by a multiple of $(m/L)(1/d)$, where $d$ is the dimension, and $L/m$ is the conditional number of the negative log of the target density.
Section F provides an alternative quantitative relationship between $\|T\|_{\Pi}$ and $\|\hat{T}\|_{\Pi}$, where $T$ and $\hat{T}$ are the exact and hybrid random-scan Gibbs kernels defined in (1) and (2) of the main text. The relationship is based on arguments from [64]. It is then compared to the quantitative relation given in Corollary 4 of the main text.
@@ Appendix E A proximal sampler
This section contains an application of Proposition 7 from the main text, taken from [2]. This application, which involves hybrid proximal samplers, may be interesting in its own right.
@@
In this case, the data augmentation chain associated with the Markov transition kernel (Mtk) $S$, where $S(y,\mathrm{d}y^{\prime})=\int_{\mathbb{R}^{d}}\Pi_{2,y}(\mathrm{d}z)\,\Pi_{1,z}(\mathrm{d}y^{\prime})$ is called a proximal sampler [43]. It can be particularly useful when $M_{1}$ is ill-conditioned, although later on in our analysis of its hybrid version, we focus on the case where $M_{1}$ is well-conditioned for tractability. This sampler was studied by Lee, Shen and Tian [44] and Chen et al. [21]. In particular, the following holds.
@@
[21, Theorem 4] Suppose that $m>0$. Then $\|S\|_{M_{1}}\leq 1/(1+m\eta)$.
@@
Proposition 7 and Corollary 9 (both from the main text) provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{1,z}\|_{\Pi_{1,z}}$. In particular, combining the bound in Proposition 7 of the main text with Lemma 30 yields the following.
@@ Appendix F A bound from Roberts and Rosenthal (1997)
As mentioned in Section 1 of the main manuscript, [64] is a pioneering work on the convergence properties of hybrid Gibbs samplers. Roberts and Rosenthal [64] did not provide explicit bounds relating $\|T\|_{\Pi}$ to $\|\hat{T}\|_{\Pi}$, where $T$ and $\hat{T}$ are the exact and hybrid random-scan Gibbs kernels defined in (1) and (2) of the main text. However, it is possible to use their argument to obtain a quantitative relationship. In this section, we derive such a relation, and compare it to Corollary 4 of the main text.
We first review the relevant result in [64]. We adopt the notations from Section 2.1 of the main text. For $i=1,\dots,n$, let $\Pi_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be such that $\Pi_{i}f(x)=\int_{\mathsf{X}}\Pi_{i,x_{-i}}(\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x_{-i}^{\prime})\,f(x^{\prime})$, and let $Q_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be such that $Q_{i}f(x)=\int_{\mathsf{X}}Q_{i,x_{-i}}(x_{i},\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x_{-i}^{\prime})\,f(x^{\prime})$. Then, $T$ and $\hat{T}$, as defined in (1) and (2) of the main text, can be written as $T=\sum_{i=1}^{n}p_{i}\Pi_{i}$ and $\hat{T}=\sum_{i=1}^{n}p_{i}Q_{i}$, respectively.
@@
Under the assumptions in Proposition 33, one may find $t$ sufficiently large so that
$n^{-t}\sum_{i=1}^{n}C_{i,t}<n^{-(t-1)}(1-\|T\|_{\Pi}).$
@@
The quantity $C_{i,t}$ is slightly mysterious. Using the notation $f_{i,x_{-i}}(x_{i})=f(x)$ in the proof of Theorem 2 from the main text, we may write it as follows:
@@
Assume, as in Corollary 4 of the main text, that $\|Q_{i,y}\|_{\Pi_{i,y}}\leq C$ for some constant $C\in[0,1]$. Then
@@
Hence, (23) is looser than the lower bound in (12) from Corollary 4 of the main manuscript. For a rough quantitative comparison, fix $T$ and assume that $C=1-\varepsilon$ where $\varepsilon=o(1)$. Take $t=\lceil-\varepsilon^{-1}\log[(1-\|T\|_{\Pi})/2]+O(1)$. Then
@@
where $\Theta$ means the left and righ hand sides are of the same order. Then the lower bound in (23) is exponential in $-1/\varepsilon$, while the lower bound in (12) of the main text is proportional to $\varepsilon$. In the example in Section 6.1.2 of the main text, $\varepsilon$ is inversely proportional to a polynomial of some dimension parameters. In that example, if $k=n$, $m_{i,y}/L_{i,y}$ is lower bounded by a positive constant $1/\kappa$, and $d_{i}=d$ for $i=1,\dots,n$, then one may take $\varepsilon=C_{0}/(\kappa d)$. In this case, (12) gives the bound $1-\|\hat{T}\|_{\Pi}\geq C_{0}/(\kappa d)(1-\|T\|_{\Pi})$, while (23) gives $1-\|\hat{T}\|_{\Pi}\geq c[(1-\|T\|_{\Pi})/2]^{(\log n)\kappa d/C_{0}}$ when $d$ is large, where $c$ is a positive constant. Evidently, the former is much tighter.